A motorist drives south at 20 m/s for 3 min, then turns west and travels at 25 m/s for 2 min, and finally travels northwest at 30 m/s for 1 min. For this 6 min trip find the average velocity.
1
Expert's answer
2020-06-01T14:27:51-0400

Let the motorist starts his journey from the origin "A=(0,0)" and reach to points "B,C,D" as given in the problem.

Now,consider the velocity in south,west and north-west direction are "v_s.v_w \& v_{nw}" respectively and also respective travel times are "t_s,t_w,t_{nw}" .


Thus from the figure, clearly,

"u=v_st_s=3600m"

"v=v_wt_w=3000m\\
w=v_{nw}t_{nw}=1800m"

Thus from the figure we can clearly see the final displacement vector is

"\overrightarrow{r}=(-1800\cos(45^{\circ})-3000,-1800\sin(45^{\circ})-3600)"

and initial displacement vector is "(0,0)" .


Now,

"\overrightarrow{V_{av}}=\frac{\Delta \overrightarrow{r}}{\Delta t}\\
\implies \overrightarrow{V_{av}}= \frac{(-1800\cos(45^{\circ})-3000,1800\sin(45^{\circ})-3600)}{360}\\
\implies \overrightarrow{V_{av}}=(-11.86,-6.46)=-11.86\hat{i}-6.46\hat{j} \:m/s"

Hence, magnitude of "\overrightarrow{V_{av}}" is

"V_{av}=\sqrt{(-11.86)^2+(-6.46)^2}=13.50m/s"